OpenStudy (anonymous):

Verify the identity. 1 - sin x / cos x = cos x / 1+ sin x

OpenStudy (anonymous):

\[\frac{1-\sin x}{\cos x}=\cdots\\ \frac{1-\sin x}{\cos x}\cdot\frac{1+\sin x}{1+\sin x}=\cdots\\ \frac{1-\sin^2x}{\cos x(1+\sin x)}=\cdots\\ \frac{\cos^2x}{\cos x(1+\sin x)}=\cdots\]

OpenStudy (anonymous):

Is that it ?

OpenStudy (anonymous):

You have a factor of \(\cos x\) in both numerator and denominator, so you can remove those, leaving you with the desired right side of the equation.

OpenStudy (anonymous):

How would I do that

OpenStudy (anonymous):

Do you understand why \(\dfrac{x}{x}=1\)? You would apply this reasoning to say that \[\frac{\cos^2x}{\cos x(1+\sin x)}=\cdots\\ \frac{\cos x}{\cos x}\cdot\frac{\cos x}{1+\sin x}=\cdots\\ 1\cdot\frac{\cos x}{1+\sin x}=\cdots\\ \frac{\cos x}{1+\sin x}=\cdots\]

OpenStudy (anonymous):

That would equal 1 + cotx right ? If That does then it would be 1 +cot x = cscx

OpenStudy (anonymous):

No. There are no cotangents involved here. Also, \(1+\cot x\not=\csc x\). You're thinking of \(1+\cot^2x=\csc^2x\).

OpenStudy (anonymous):

I dont follow I am sorry can u just tell me the answer but explain it step by step please ?

OpenStudy (anonymous):

You start off with \[\frac{1-\sin x}{\cos x}=\frac{\cos x}{1+\sin x}\] Notice the factor of \(1+\sin x\) in the denominator on the right side. That should be a clue as to what to do for the first step. How do you get \(1+\sin x\) in the denominator on the left side? Multiply the \(\cos x\) on the left by \(1+\sin x\). However, whatever you do the denominator must be done to the numerator to maintain equality. So, you multiply the left side by \(\dfrac{1+\sin x}{1+\sin x}\): \[\frac{1+\sin x}{1+\sin x}\cdot\frac{1-\sin x}{\cos x}=\frac{\cos x}{1+\sin x}\] This doesn't change the value of the left side, since (anything)/(same anything) = 1, and multiplying anything by 1 does not change the value. Make sense so far?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Now we distribute: \[\frac{1+\sin x}{1+\sin x}\cdot\frac{1-\sin x}{\cos x}=\frac{\cos x}{1+\sin x}\] becomes \[\frac{(1+\sin x)(1-\sin x)}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\\ \frac{1-\sin^2x}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\] The numerator is a difference of squares. \((a+b)(a-b)=a^2-b^2.\) Next, we use the following Pythagorean identity, \(\sin^2x+\cos^2x=1\). Rearranging a bit and solving for \(1-\sin^2x\), we have \(1-\sin^2x=\cos^2x\). So, we substitute: \[\frac{\color{red}{1-\sin^2x}}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\] becomes \[\frac{\color{red}{\cos^2x}}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\] Are you still with me?

OpenStudy (anonymous):

\[\frac{\cos^2x}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\] is the same as \[\frac{\cos x\cos x}{\cos x(1+\sin x)}=\frac{\cos x}{1+\sin x}\] or \[\frac{\cos x}{\cos x}\cdot\frac{\cos x}{1+\sin x}=\frac{\cos x}{1+\sin x}\] And since \(\dfrac{\cos x}{\cos x}=1\), we are left with \[\frac{\cos x}{1+\sin x}=\frac{\cos x}{1+\sin x}\] Thus the identity (equality) is verified, and we're done.

OpenStudy (anonymous):

Ok thanks so much I think I have one more question if you could still help me that would be great if not thats fine

OpenStudy (anonymous):

Sure, would you mind posting it as another question? In case I can't help, someone else might be able to.

OpenStudy (anonymous):

Yeah I will I am just writing down everything for the problem

OpenStudy (anonymous):

Okay, and you're welcome!

OpenStudy (anonymous):

Thanks again